OpenStudy (anonymous):

True or False. The equation 4x2 + 16 = 0 is an example of a quadratic equation. Explain. I think it's false because a quadratic equation needs to be in the form of ax^2 + bx +c = 0... I have a feeling I'm wrong though..

OpenStudy (zarkon):

it is of that form...a=4,b=0,c=16

OpenStudy (unklerhaukus):

yeah b is just zero, and that is fine

OpenStudy (mertsj):

It contains a second power of x as its highest power so it is quadratic.

OpenStudy (mertsj):

In other words, it is a second degree equation so it is quadratic.

OpenStudy (anonymous):

but could C be zero too? ax^2 + 16 + 0 =0 ?

OpenStudy (anonymous):

ooooh ok, the second degree equation rings a bell

OpenStudy (mertsj):

Yes a quadratic equation is of the form ax^2 +bx+c where a is not 0. b and c can be 0 but not a

OpenStudy (mertsj):

x^2=0 x^+3x=0 x^2-5x+7=0 All are quadratic equations.

OpenStudy (unklerhaukus):

*x^2+3x=0